OpenStudy (anonymous):

(x+1)^2=25 solve plzzz

OpenStudy (anonymous):

take sqrt on both sides (x+1) = 5 or -5 Case1 x+1=5 so x=4 Case2 x+1=-5 so x=-6

OpenStudy (anonymous):

the closest i got was x=10,13 but i dont think i did it right at all

OpenStudy (anonymous):

ty